OpenStudy (anonymous):

Need help w/ 8th grade math!!! (:O Here's the problem: Loon lake has a killer plant problem. Currently 5000 square feet is covered with the plant. the area is growing with a factor of 1.5 each year. i have to make a table to show the area covered with the plant for the next 5 years but i dont understand how to do problem!! or get the area covered each year!! help please!!

OpenStudy (anonymous):

First make an equation. Let's say, L = area loon lake's plant covered each year y = number of years.

OpenStudy (anonymous):

L = 5,000 (because they had 5,000 square feet already covered before hand) + x1.5 (because every year it grows 1.5) L = 5000 + x1.5

OpenStudy (anonymous):

Let's try the first year together.

OpenStudy (anonymous):

It has been 1 year now. We plug in 1for x, because x is equal to the number of years. L = 5000 + (1)1.5 Now we solve! L = 5000 + 1.5 L = 5001.5 So, for the first year, 5,001.5 acres were covered.

OpenStudy (anonymous):

Can you do the rest?